OpenStudy (anonymous):

5/2- 3x= 6/7 -4x 3x +4x= 6/7 +-5/2 7x= 6/7*2 + -5/2*7 7x=12/14 + 10/14 7x/7 = 22/4 x= 5 1/2 What did I do wrong?

OpenStudy (anonymous):

is it \[\frac{5}{2}-3x = \frac{6}{7}-4x\] If so, in your first step it should be\[-3x+4x = \frac{6}{7}-\frac{5}{2}\]

OpenStudy (anonymous):

then you get \[x=\frac{2*6}{2*7} - \frac{7*5}{7*2} = \frac{12}{14} - \frac{35}{14}\]

OpenStudy (anonymous):

someone gave me this as the right answer4x-3x = 6/7-5/2 x= ((6*2)-(5*7)/14 = (12-35)/14 x = -23/14 x= -1 9/14 The answer is -1 9/14 simplest form